You have 100 coins that are only nickles and dimes. All the coins add up to 7.00. How many nickles do you have? How many dimes to you have?
Answer by checkley79(3341)


.05n+.10d=7
n+d=100
n=100-d
.05(100-d)+.10d=7
5-.05d+.10d=7
.05d=7-5
.05d=2
d=2/.05
d=40 number of dimes.
n+40=100
n=100-40
n=60 number of nickles.
PROOF:
.05*60+.10*40=7
3+4=7
7=7